# Yuang Li, Yinglu Li, Min Zhang, Chang Su, Mengxin Ren, Xiaosong Qiao, Xiaofeng Zhao, Mengyao Piao, Jiawei Yu, Xinglin Lv, Miaomiao Ma, Yanqing Zhao, Hao Yang

### v3

@@ 5 Conclusion
This paper presents CB-Whisper, an innovative ASR framework that leverages the prior knowledge of entity words to enhance their recalls. Our model consists of an OV-KWS module that exploits the hidden states of the Whisper encoder. The entities detected by this module serve as prompts for the Whisper decoder. We devise a multitask training strategy that simultaneously optimizes OV-KWS and ASR tasks. We also propose a more convenient method to apply the OV-KWS model to various Whisper models without finetuning using spoken form prompts. We evaluate our method on four test sets and show that it achieves substantial improvements in both MER and entity recall. In future work, we aim to enhance the efficiency of the OV-KWS module to enable it to handle a large keyword database and conduct large-scale experiments by distilling speech data from the original Whisper model, to boost the performance of Multitask finetuning in the general domain.

### v4

@@ 5 Conclusion
This paper presents KWS-Whisper, an innovative ASR framework that leverages the prior knowledge of entity words to enhance their recalls. Our model consists of an OV-KWS module that exploits the hidden states of the Whisper encoder. The entities detected by this module serve as prompts for the Whisper decoder. We devise a multitask training strategy that simultaneously optimizes OV-KWS and contextual-ASR tasks. We also explore more convenient methods to apply the OV-KWS model to correct ASR errors and to prompt frozen Whisper models. Evaluated on four test sets, our method is shown to achieve substantial improvements in both MER and entity recall. In future work, we plan to enhance the efficiency of the OV-KWS module and enable it to handle a large keyword database and conduct large-scale experiments by distilling speech data from the original Whisper model, thereby boosting the performance of multitask fine-tuning in the general domain.